Systems Staff Meeting Notes 7/19/99

Systems Staff Meeting Notes
July 19, 1999

Present: Brian Burton, Bill Covey, Debra Harris, Winston Harris, Bobbie Meng, Everett Phillips, Suzy Shaw, Ying Tang
Announcements
  1. Brian Burton will be leaving Systems for a new job with Marion County. He is also pushing his karate activities further, helping to open a new DoJo in Dunnellon, Florida. A new search will begin for another Programmer upon Brian's departure at the end of next week.
  2. The search for a new Network Coordinator is in progress. This week we'll be interviewing a candidate who currently works at the University Police Department. He has a good deal of network experience, especially in Administration.
  3. Christian Conrad has left the building. We still need to get his Intellikey back. He currently resides in Tigert.
  4. Everett's soccer team tied with Orange Crush, 2 to 2.
  5. Travel money control will now be given to each department head. Bill proposes to use the $1000 allocated to Systems for certification testing: A+, Linux, Novell, MS & CISCO.
  6. Enrollment in on-campus computer classes can now be taken with a fee waiver. Schedules can be adjusted as necessary.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Cow shipment arrived last week. Distribution list was issued this morning, they are ready to go. Will forward the list to Liaisons & Middle Managers. In order to pick up new machines, staff must return the old ones that are being replaced. They'll also need to verify that there is a Systems staff member available to open the vault door.
Liaisons
  1. Liaison Training will take place the week of Aug. 9 & again the week of Aug. 16.
  2. Everett & Bobbie need to attend separate sessions.
  3. Pre & post tests will be given. Only with extra study can certification be achieved.
Networking
  1. Phantom IPX problems - there have been problems with morning log-ons of public machines. Solution is to log on one at a time for now, until NERDC can rectify the problem.
  2. Novell log-in problems - Smathers needs to have the "not logged-in" cleared out daily. KillCon is not working properly.
  3. Documents desk wiring - need to check on it.
Software Projects Status
  1. Personnel - Evelyn's OPS database is not working properly. Winston notes that a modification he made is not working correctly.
  2. IR database needs input before it can move on.
  3. Winston is meeting with Alena on SQL Access.
  4. Trouble Ticket Viewer problems have been resolved. It is now ready for distribution.
  5. PPD deal is probably dead.
DP Coordinator
  1. Student upheaval - Brian Monk is no longer working for us. Leo is on probation. New student (a friend of Lindsay's) will be hired. Her name is Runa Diwadkar.
  2. Compustat - on hold pending Bill/Bevis connection.
Calendar
  1. Rich out July 19-30.
  2. Debra out Aug. 16-20.
  3. Bill will soon be out for a week (at least) on a southeastern tour of University Systems departments.


Back to the Systems Homepage

Last updated February 18, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu